### v2

# Just-in-time and distributed task representations in language models

Abstract
Many of language models’ impressive capabilities originate from their in-context learning: based on instructions or examples, they can infer and perform new tasks without weight updates. In this work, we investigate when representations for new tasks are formed in language models, and how these representations change over the course of context. We focus on “transferrable” task representations—vector representations that can restore task contexts in another instance of the model, even without the full prompt. We show that these representations evolve in non-monotonic and sporadic ways, and are distinct from a more inert representation of high-level task categories that persists throughout the context. Specifically, when more examples are provided in the context, transferrable task representations successfully condense evidence. This allows better transfer of task contexts and aligns well with the performance improvement. However, this evidence accrual process exhibits strong locality along the sequence dimension, coming online only at certain tokens—despite task identity being reliably decodable throughout the context. Moreover, these local but transferrable task representations tend to capture minimal “task scopes”, such as a semantically-independent subtask. For longer and composite tasks, models rely on more temporally-distributed representations. This two-fold locality (temporal and semantic) underscores a kind of just-in-time computational process that language models use to perform new tasks on the fly.
1 Introduction
Much of the excitement about large language models began with the discovery that they exhibit In-Context Learning (ICL; Brown et al.,, 2020): the emergent ability to learn tasks from few-shot examples in context. This discovery has led to a variety of works exploring the behavioral features of ICL (e.g. Sclar et al.,, 2024; Min et al.,, 2022). Other works have studied the dynamics of ICL, and how performance improves with increasing numbers of few-shot examples (Agarwal et al.,, 2024; Anil et al.,, 2024). The strong behavioral success of ICL led to substantial interest in understanding the mechanistic basis of these capabilities, leading to discoveries such as induction heads (e.g. Olsson et al.,, 2022) and how ICL implicitly refines a model of in-context evidence (e.g. Akyürek et al.,, 2022; Von Oswald et al.,, 2023).
Recently, several works have identified internal, vector-form task representations that can be extracted from a model’s forward pass on a few-shot prompt (Todd et al.,, 2024; Hendel et al.,, 2023). These task representations not only capture general task information, but can be used to restore the appropriate task context during the model’s forward pass on a zero-shot prompt. This transfer effect is observed by intervening with that representation at the appropriate place in the model’s residual stream. Such intervention reinstates the task context and allows the model to perform the task without any explicit instructions or demonstrations. These “transferrable” task representations have been shown to exist across a variety of tasks and presentation formats, and even capture transferrable task knowledge across modalities (Davidson et al.,, 2025; Huang et al.,, 2024; Luo et al.,, 2024).
The discovery of transferrable task representations raises several intriguing questions about their dynamics and generality. How and when are these internal task representations formed throughout the context? How do these dynamics depend upon task complexity? A simple, intuitive hypothesis is that these representations develop gradually, e.g. at a rate that depends on task complexity. The representations might accrue evidence across tokens and refine monotonically into a more stable and robust task representation. This view aligns with the behavioral findings that models perform better with more examples in-context (Agarwal et al.,, 2024; Anil et al.,, 2024).
We set out to investigate these transferrable task representations to understand how the dynamics of ICL are exhibited within them. Our findings suggest a more complicated picture of the computational process that support language models’ adaptation to a new task:
  • We find two types of task representations in language models: an inert representation of task identity is more continuously present throughout the context, but transferrable task representations only activate sporadically at key tokens.
  • The fleeting but transferrable task representations condense evidence from multiple examples. However, they also tend to capture minimal task scopes, as their ability to guide model behavior decays over longer generation and across independent subtask contexts.
  • Models do not appear to condense task knowledge into local task representations in more complicated tasks that require state tracking or chaining multiple subtasks together.
  • Finally, models can form distinct representations to support generating the same responses when solving tasks independently vs. as part of a broader context.
Overall, these results give us a window into language models’ changing state when inferring and solving new tasks in context, but paint a complex and nuanced picture of the dynamics of this state. There are different types of task representations—identifiable vs. transferrable—that evolve over the context in distinct ways. The representations of tasks also depend on the task complexity and the surrounding context structure in which a task is embedded. These results may have implications for both the science of understanding models, and practical applications of mechanistic interpretability for analysis and safety.
Figure 1: Understanding how task representations develop over context. A. A schematic of extracting transferrable task representations and restoring task contexts in zero-shot settings. The highlighted tokens indicate the source and target for extracting and injecting task representations. B. Transferrable task representations restore task accuracy on zero-shot prompts. Results are aggregated over all models for simple tasks (see Appendix A). Error bars indicate the 95% CI over tasks. C. An overview of the development of different task representations over context. Solid bars: recontextualized zero-shot accuracy for task vectors extracted from different tokens. Transparent bars: task identity decoding accuracy from different token representations.
2 Methods
Tasks
For our analyses, we built upon tasks from prior work on transferrable task representations (Hendel et al.,, 2023; Todd et al.,, 2024). The tasks we examine include a diverse set of natural language tasks (e.g., finding the antonym of a query word or translating an English word to French) and algorithmic tasks (e.g., counting or extracting a target word from a list of input words). In addition to these simple, single-token generation tasks from the previous literature, we also test a range of new tasks to explore model behavior in longer generation settings. These include: repeating a simple task three times (e.g. antonym x 3 requires finding the antonyms of three input words), extracting multiple words from a query word list (e.g., choose both the first and the last word in the list), and reversing or shifting an entire word list. Finally, we also explore a set of “mixed-generation” tasks, where the model needs to infer and perform different tasks on each input item. See Appendix A for the full set of tasks.
There are 512 query-answer pairs for each task (except for two smaller datasets: country-capital contains 197 samples, and product-company contains 494 samples). These query-answer pairs are formatted into few-shot prompts with alternating “Q:” and “A:” turns, as shown in Figure 1A.
Models
In the main paper, we present results on the open-weight pre-trained 4B, 12B, and 27B Gemma V3 models (Team et al.,, 2025). In Appendix C, we also show that the main findings replicate on the 4B, 8B, and 14B Qwen3 models (Yang et al., 2025a, ).
Extracting transferrable task representations
We primarily investigate task vectors discovered in Hendel et al., (2023) as a window to study language models’ transferrable task representations. We extract task vectors from few-shot prompts consisting of query-answer pairs and a test query, as shown in Figure 1A. Task vectors are the layer residual activations extracted from the last token before answer generation (in the example in Figure 1A, this corresponds to the highlighted colon token). Hendel et al., (2023) showed that task vectors can reinstate task performance on a different query even without any prior context. Specifically, when task vectors are patched onto (i.e., overwrite) the layer residual activations of the last token, they can recontextualize the model with the appropriate task context and enable the model to generate the task output without any prior few-shot examples.
We replicate and extend the procedure outlined in Hendel et al., (2023). For each model and task, we first search for the layer that best captures the task representation, using 50 queries from the dataset as the development set and in an 8-shot setting. As in prior work, we replace the real test queries with dummy queries sampled from the dataset to extract query-agnostic, general task representations. We searched among every 3 layers starting at layer 2 (0-indexed) for the 4B and 12B models (covering both the local-attention layers and global attention layers in Gemma V3 models; Team et al.,, 2025), and every 6 layers starting from layer 5 in the 27B model (covering the global attention layers). The layer that restores the highest task accuracy on zero-shot prompts in the development set is designated as the layer that best captures the representation for a given task. This best layer is subsequently used to extract task vectors and restore task contexts for the remainder, held-out queries in the dataset. Consistent with prior results, we generally find that task vectors extracted and injected at middle layers restore the highest task accuracy on zero-shot prompts, for all model sizes.
Evaluating task transfer
We compare the average accuracy of the sampled responses across three settings: standard zero-shot, recontextualized zero-shot (with task vector intervention), and few-shot (with examples in context). For simplicity, responses for all tasks are graded by exact string matches against the ground-truth answer. This underestimates the model performance in some tasks (e.g. for antonym and translation tasks), but we use the same grading scheme across all settings and compare relative performances. For longer-generation and mixed-generation tasks, we evaluate each of the multiple outputs separately by exact match (e.g., in antonym x 3, we compare each of the three output words with the correct answer), and report the average accuracy across all output units.
Examining the dynamics of transferrable task representations
Once we determine the best layer for each task using the last colon token, we evaluate how well the colon token representations condense information from multiple examples in a prompt. We do this by extracting task vectors at the colon token in different k-shot prompts and patching onto zero-shot prompts, then comparing the recontextualized zero-shot accuracy. We repeat this analysis for k in 0, 1, 2, 4, 8, 16, 32. In an earlier experiment on a subset of the tasks, we also experimented with allowing the best layer to vary depending on k, but found very similar results overall. We show these layer search results across different k’s in Figure 2A, but otherwise focus on results from reusing the best layer from the layer search on 8-shot prompts for other k’s.
We then studied whether transferrable task representations form in tokens other than the colon token. We extracted layer residual token activations for other format tokens in the context, including the “Q”, the “:” following “Q”, the “A”, and the new-line token before “A”. We patched these token activations onto the corresponding token in the zero-shot prompt at the same layer. For each of the non-colon tokens, we repeat the search for the layer that best captures task representations. All token representations are evaluated on the extent to which they restore task accuracy on zero-shot prompts.
Decoding task identity
We also studied whether token representations throughout the context contain robust task identity information. We trained simple linear decoders to predict the task category from the layer residual activations of different tokens. At each layer and token combination, we used 100 token instances for each task to train and test a task identity decoder. All decoders were trained for 20 epochs, with a batch size of 256 and a learning rate of 0.01. We report the decoding accuracy across the 25% held-out test representations.
3 Results
How do language models infer new tasks in-context? We leverage the transferability of task contexts to understand how models accrue evidence and refine task representations. We find that language models indeed form stronger task representations that aggregate in-context evidence. However, this evidence accrual process is non-gradual and happens in a sporadic way. In particular, models only form effective task representations that transfer task contexts at certain tokens (Figure 1C, also see Figure 2A). This is in strong contrast to a persistent sensitivity of task identity across tokens in the context (Figure 2B). We also find that the surprisingly local transferrable task representations tend to only capture a minimal “task scope.” Below, we discuss these findings in more detail.
Figure 2: Transferrable task representations activate sporadically at key tokens, but task identity representations persist throughout the context. A. Recontextualization accuracy when each token representation is used to restore task contexts in zero-shot settings. B. Task identity decoding accuracy (among 14 tasks) for token representations at different layers and positions. This figure plots aligned sequences across different samples and tasks; since exact positions differ depending on the sample, the indices shown in the labels are approximate. See results for other models in Figure S4.
Figure 3: Sporadic & inconsistent evidence accrual in language models. A. Task vectors extracted from the last colon token in each example capture evidence accrual on most tasks (12 out of 14). However, on two “hard-to-transfer” tasks, task vectors do not capture this evidence accumulation, even though the models (behaviorally) do learn from more examples. The solid bars indicate recontextualized zero-shot accuracy (via task vectors), and light bars in the background indicate few-shot accuracy (without task vectors). The dotted lines indicate the ratio of the recontextualized zero-shot accuracy against few-shot accuracy. B. Most other format tokens in the context do not robustly form transferrable task representations that support recontextualization on zero-shot, but task identity is reliably decodable in their residual activations. Here, we report the task identity decoding accuracy at the mode best layer at which transferrable task representations form in the second “:” token. See the main text for more details.
3.1 Local task representations can accrue evidence
Transferrable task representations reflect evidence accrual.
Consistent with the behavioral gain from including more examples in-context (e.g., Anil et al.,, 2024), we find that transferrable task representations also reflect increased task certainty with increased context (Figure 3A). Task vectors extracted following more examples are better at restoring task performance in zero-shot settings, such that the ratio between the recontextualized zero-shot accuracy and few-shot accuracy stays relatively stable across the number of examples. This suggests that language models condense information from multiple examples and form better task representations, even though the task representations extracted are fairly local (i.e., from a single token in the few-shot prompts).
… but not for all types of tasks.
However, we did not observe strong evidence accrual in 2 out of the 14 simple tasks (Figure 3A, hard-to-transfer tasks). These two tasks are count_color_in_3 and count_fruit_in_3. For all three model sizes, task vectors extracted from 32-shot prompts recovered below 50% of the few-shot accuracy with 32 examples in the context. In other words, local task representations for these tasks were not able to take advantage of more examples for task transfer to zero-shot settings, even though models improved substantially at solving the task when given more examples in the prompt. Interestingly, the same counting tasks were hard-to-transfer for Qwen3 models as well (see Figure S7A). One possibility is that these tasks require more state-tracking, which may necessitate additional inference processes that the models do not condense into local task representations. Alternatively, these inference processes cannot be effectively re-activated by the injection of the extracted task representations.
How evidence accrual leads transferrable task representations to converge.
As models appear to successfully accrue evidence in these highly local representations, we sought to understand how the task representations themselves changed over more examples (Figure 4A). For this analysis, we look at the task representations extracted from the mode best layer across different tasks. This is to control for magnitude differences of the residual activations across layers and make a fair comparison. Although the best layer for transferrable task representations sometimes differ across tasks, the best layers tend to reside in the middle layer range across all model sizes, consistent with prior findings (Hendel et al.,, 2023).
In Gemma V3 models, as we increase the number of examples in context, we generally found reduced variance among task vectors extracted from different k-shot prompts. This can signal that in-context task representations tend to denoise or converge to more stable representations as models gain evidence. The magnitude (L2-norm) of the task vectors also tends to decrease over time. However, for both the variance and magnitude, there were considerable differences between tasks and model sizes. Some tasks seem to converge to stable representations faster (i.e., with fewer examples; see also a visualization of the representational trajectories in Figure S2). For certain tasks, the magnitude of the task representations first increases then decreases given more examples. We also note that both the variance and magnitude of task vectors tend to increase per more examples in some Qwen3 models (Figure S8), suggesting that different models may develop different strategies for refining task representations.
Figure 4: Analyses of extracted task representations in Gemma V3 models. A. The extracted task vectors (at the last colon token) tend to decrease in both variance and magnitude with more examples, exhibiting a general tendency to condense evidence and converge onto stable task representations. The solid line shows the average across tasks. The transparent lines show the individual tasks. B. Extracted task vectors from the 27B model form distinct clusters. The numbers label the centroid for each task (see legend and results for other models in Figure S5). Task vectors are similar but distinguishable when a task is evaluated independently vs. embedded within a larger task structure. For example, representations for antonym (0), antonym x 3 (14), and where antonym appears as a first task in a mixed-generation task chain (24&27) are close but distinct.
3.2 Task representations exhibit temporal locality
The analyses above confirm that transferrable task representations benefit from increasing in-context evidence and converge to better representations. To understand the full temporal profile of this accrual process, we repeated the task vector recontextualization experiment on other tokens in the prompt, which revealed that the transferrable task representations do not strengthen monotonically.
Transferrable task representations are not found in most tokens.
We tested whether extracted representations from other format tokens in the prompt can also restore the corresponding task contexts. These include “Q”, the “:” following “Q”, “A”, and the new-line token before “A”, which are all shared across examples, tasks, and contexts. As before, we patched the activations at the same layer, but onto the corresponding format token instead of the last colon token in zero-shot prompts. As shown in Figure 3B, transferrable task representations generally do not form in the residual activations across layers in these tokens. This is true across the number of examples provided in the prompt, leading to the developmental trajectory of transferrable task representations shown in Figure 1C.
We observed nearly zero recontextualized zero-shot accuracy for all these tokens in most tasks, except some restoration success in product-company, color_v_animal_3, choose_first_of_5, and the longer-generation tasks discussed below (see Figure S3). In Qwen3 models, there also appear to be some restoration successes with representations extracted from the “Q” token (see Figure S7B and Figure S11). These partial successes are likely driven by the fact that the first answer token sometimes match the first input token across a subset of the tasks. In general, it seems that an effective, transferrable task representation in language models only forms sparingly; in few-shot settings, this often means a just-in-time task representation at the token before answer generation.
… but a robust task identity signal persists throughout the context.
Intriguingly, however, task identity is almost perfectly decodable in the representations extracted from all the different format tokens, even though the formats are shared across all tasks. We report the decoder accuracy from the layer with the best transferrable task representations in Figure 3B, but found high task identity decoding accuracy across most tokens throughout the context (Figure 2B). The decodability success may be partly due to vocabulary differences between some tasks, but we show that decoding accuracy remains high even for a restricted subset of tasks with shared vocabulary (Figure S4C). Both task identity and task transferability do not occur until at least one full example is presented, but accurate task identity representations form much earlier than transferrable task representations (Figure 1C). This suggests that the model is generally task-sensitive, but instantiates transferrable task representations only at particular timepoints in the context.
3.3 Task representations exhibit scope locality
We have seen evidence that transferrable task representations tend to be temporally local. That leads to the question of whether they have a lasting effect over generation. That is, are the restored task contexts in the zero-shot forward pass also fleeting in nature? Prior work has mostly focused on simple, single-token output tasks. To study the scope of task representations, we tested to what extent restored task contexts can support longer generation beyond the first token. Building on the simple tasks from prior work, we evaluated models on a set of longer-generation and mixed-generation tasks, including repeating a simple task multiple times on different input words, list-level tasks that operate over multiple words, and inferring/performing different tasks on different words (see Methods and Appendix A).
Figure 5: Reinstantiated task contexts in longer- and mixed-generation tasks often decay over generation, especially for tasks that can be decomposed into semantically-independent subtasks. This suggests a tendency for models to only activate transferrable representations for small task scopes. A. Bar plot: recontextualized zero-shot accuracy compared to zero-shot and 8-shot accuracy on longer-generation tasks. Line plots: recontextualization accuracy for each output unit across models, conditioned on sequences where models generated full correct responses with eight examples in-context. An output unit usually corresponds to a single word and is occasionally a short phrase (e.g. the capital of a country). B. Visualization as in A, but for mixed-generation tasks.
Transferrable task representations tend to support a semantically-independent task scope.
In these experiments, we find further evidence of the locality of transferrable task representations. Overall, the recontextualized zero-shot accuracy of tasks that require longer and mixed answers is substantially lower than that in tasks that require shorter answers (Figure 5, bar plots; also see Figure S1). Across a range of tasks, we find that the recontextualized zero-shot accuracy decreases for each output word (Figure 5, line plots; see similar results on Qwen3 models in Figure S13), suggesting that the restored task contexts “fade” over longer generation.
Models seem to be consistently decomposing tasks and forming local task representations that capture a minimal “task scope”. For example, in choose_first_middle_last_of_5, the reinstantiated task context only supported generating the first word, suggesting that further token generations rely on additional representations instantiated by a separate mechanism while processing the previously-generated tokens. This effect is very pronounced in the mixed-generation tasks. In these cases, all models form strong local task contexts that only encapsulate the first subtask in a multi-task chain, and defer representing later subtasks. Interestingly, the extracted task representations for the same simple task when it appears independently or as a first task in a repeated or mixed-task context can be distinct (Figure 4B), even though they support generating the same responses. This can potentially reflect an inert sensitivity to the broader mixed-task context based on vocabulary differences. Additionally, these differences may also convey information relevant for later subtasks to be activated, even though they are not sufficient to directly restore execution of later subtasks.
The restored task contexts do support longer generation to some extent in many tasks, including repeating a simple task multiple times and some list-level tasks (i.e., reversing or shifting a word list). Overall, these results suggest that language models may be automatically segmenting semantically-independent task scopes when possible, such that task representations for longer or composite tasks are offloaded onto multiple tokens.
4 Related work
Since the discovery that large language models exhibit emergent in-context learning (Brown et al.,, 2020), there has been substantial interest in investigating this capability and its mechanistic basis.
From a behavioral perspective, many subsequent works have explored how ICL could develop from implicit meta-learning of data properties (Xie et al.,, 2022; Chan et al.,, 2022), and how this may relate to the broader set of language model capabilities (Chen et al.,, 2024; Lampinen et al.,, 2024). Some of this work has focused on the surprising fragility of ICL to subtle prompt changes (e.g. Sclar et al.,, 2024); conversely, others have highlighted how ICL may be overly robust, allowing “learning” common tasks even if the labels are randomized (Min et al.,, 2022). One particularly relevant focus of behavioral work on ICL has been on the dynamics of in-context learning; for example, how adding many example shots can improve performance on difficult tasks, or even those discouraged in post-training (Agarwal et al.,, 2024; Anil et al.,, 2024).
From a mechanistic perspective, Olsson et al., (2022) showed how induction heads could support in-context learning, and other work has studied how they might develop over training (Edelman et al.,, 2024; Singh et al.,, 2025). More recently, attention has turned to cases where models may create internal task representations. These are representations that can be extracted from few-shot prompts and then injected (without the few-shot examples in context) to induce task performance. Hendel et al., (2023) demonstrated an instance of such task representations: representations at intermediate layers of the model at a key token that can be injected to mimic the effect of a few-shot prompt. Concurrent work from Todd et al., (2024) identified “function vectors,” which aggregate the effects of multiple attention heads that convey task information. Subsequent work has generalized and extended these findings, for example exploring how function vectors can emerge from instructions as well as examples (Davidson et al.,, 2025) and how task vectors capture task representations in multimodal models (Huang et al.,, 2024; Luo et al.,, 2024). Other works have explored how these representations and structures emerge over training (Yang et al., 2025b, ; Yin and Steinhardt,, 2025), and extended existing methods to more robustly restore task contexts in zero-shot settings (Li et al.,, 2024). Our work builds on these findings and use transferrable task contexts as a window to study the dynamics of in-context task representations in language models.
5 Discussion
We sought to understand the dynamics of in-context task representations that support language models’ successful learning of new tasks. Building on prior methods, we evaluated when in the context we can extract transferrable task representations from few-shot settings that restore task contexts in zero-shot settings. Our results show that in many tasks, models refine task representations over more evidence such that the representations more successfully restore task contexts. However, these transferrable task representations only sporadically activate, and seem to best support a minimal task scope (e.g., a first irreducible subtask). The dynamics of these effective, transferrable task representations strongly contrast with a general sensitivity to high-level task differences that persists throughout the context. In addition, we find many cases where models do not appear to condense global task representations into a local, token-level representation, such as tasks involving more state tracking and tasks chaining different types of subtasks together.
In general, our results complicate the intuitive picture that the computational process underlying language models’ in-context learning and task inference is smooth and gradual. Models do not appear to refine task representations at a per-token basis, not even in a step-wise manner—even on a per-example basis, the task state is not sustained, but often fades and reactivates across different tokens. That language models elect to condense task evidence onto a single token at intermediate layers rather than relying on repeated cross-token attention to aggregate information may suggest a general inductive bias to compress knowledge to local representations when possible. This may also relate to the success many works have observed on extracting transferrable task representations in broad settings, such as following instructions or images (Davidson et al.,, 2025; Huang et al.,, 2024; Luo et al.,, 2024), or capturing information for multiple possible task outcomes (Xiong et al.,, 2024).
However, one trend that arose from these investigations is that language models do not seem able to condense task information into local representations in all cases. Rather, they exhibit a tendency to form sharp local contexts for small task units, and offload broader task contexts such as mixed or multiple tasks across time. As discussed earlier, even for simple tasks where some intermediate-state tracking may be required, successful inference may need to rely on cross-token and cross-layer computation (e.g., as shown in Ameisen et al.,, 2025). In these cases, the effective restoration of the computation process may also require intervening multiple components during the forward pass. It is also possible that by overriding token activations at intermediate layers with task vectors, the models have lost any task state information formed in earlier layers. Other methods that restore task states through additive injection rather than activation patching may therefore be more successful at restoring model task states in these tasks (Todd et al.,, 2024; Li et al.,, 2024). We also note that different models may differ in the extent to which they form local or distributed effective representations across the context or layers.
An intriguing direction for future work is to study the mechanistic bases for the strong temporal and scope locality we observed in models’ in-context task representations. One possibility may be that the residual stream is more stable and easier to learn from during training. This may encourage the model to rely more on the residual stream to condense contextual task representations rather than relying on the more expensive attention operations. These learning dynamics may drive models to conform with an implicit normative consideration to not instantiate task contexts until needed and instantiate just the right scope to avoid capacity waste. Some of these features may even be exclusive to models pre-trained on natural languages (e.g. Yang et al., 2025b, ).
Limitation
We note a few important limitations of our work. First, we primarily rely on task vectors as the method to extract transferrable task representations. This means that our conclusions and speculations are bounded by the effectiveness of this method. As we discussed earlier, representations for some task contexts may be more distributed, either across tokens and/or across model layers. It would be important to confirm if similar dynamics are observed in less-constrained methods such as a multi-layer recontextualization (Li et al.,, 2024). Second, we mostly explored relatively simple tasks, including when we investigated longer-generation tasks. It’s possible that many of the dynamics we observe here would not generalize to settings with naturalistic languages, especially when the tasks are not so cleanly decomposable and a single, semantically-independent task unit is hard to define.
Conclusion
We investigated how the dynamics of in-context learning are reflected in the development of language models’ internal task representations. Our results suggest that language models do not smoothly refine a global task state in-context. While general task sensitivity persists throughout context, models appear to construct effective task representations in a “just-in-time” fashion to solve a simple task scope immediately ahead. The fleeting, minimally-scoped nature of these in-context task representations provides new insight into the models’ state of inferring and performing tasks based on new evidence.
Acknowledgments and Disclosure of Funding
We thank Yasaman Bahri, Katherine Hermann, Mike Mozer, Sjoerd van Steenkiste, and Murray Shanahan for helpful comments and support on this work.
Appendix A Tasks
Table 1: Simple/shorter-answer tasks. See Todd et al., (2024) for more details for the first nine tasks.
Task Name Example
antonym Q: true A: false
country-capital Q: Germany A: Berlin
english-french Q: queens A: reines
product-company Q: Windows XP A: Microsoft
color_v_animal_3 Q: blue, dolphin, swan A: blue
fruit_v_animal_3 Q: lime, parrot, buffalo A: lime
choose_first_of_5 Q: envelope, pasta, cake, toucan, create A: envelope
choose_middle_of_5 Q: candy, charismatic, laptop, realize, eel A: laptop
choose_last_of_5 Q: affable, believe, carefree, zoom, moray A: moray
word_length Q: negotiate A: 9
count_color_in_3 Q: snake, gold, indigo A: two
count_fruit_in_3 Q: lime, newt, bunny A: one
position_of_color_in_3 Q: monkey, oryx, white A: third
position_of_fruit_in_3 Q: pear, coyote, capybara A: first
Table 2: Longer-generation tasks.
Task Name Example
antonym x 3 Q: fall, everybody, intact A: rise, nobody, broken
product-company x 3 Q: iWork, Windows NT 3.5, OS X Yosemite A: Apple, Microsoft, Apple
all_but_color_in_3 Q: cat, black, pelican A: cat, pelican
all_but_fruit_in_3 Q: grape, butterfly, llama A: butterfly, llama
choose_first_last_of_3 Q: white, house, wallet A: white, wallet
choose_first_middle_last_of_5 Q: dolphin, beyond, curtain, pillow, intuitive A: dolphin, curtain, intuitive
reverse_all_of_3 Q: donut, sad, who A: who, sad, donut
reverse_all_of_5 Q: she, honest, out, test, frog A: frog, test, out, honest, she
shift_all_of_3 Q: piano, cougar, jackfruit A: cougar, jackfruit, piano
shift_all_of_5 Q: agreeable, flamingo, short, around, jovial A: flamingo, short, around, jovial, agreeable
Table 3: Mixed-generation tasks.
Task Name Example
antonym + product-company Q: opponent, iDisk A: ally, Apple
english-french + antonym Q: liberal, continue A: libéral, stop
product-company + english-french Q: Alfa Romeo MiTo, mask A: Fiat, masque
antonym + country-capital + english-french Q: upper, Greece, artists A: lower, Athens, artistes
english-french + product-company + antonym Q: system, Lancia Flavia, unlucky A: système, Fiat, lucky
country-capital + antonym + product-company Q: Gambia, heavy, Game & Watch A: Banjul, light, Nintendo
Appendix B Additional Gemma3 results
Figure S1: Recontextualization accuracy for all tasks. Task vectors extracted from 8-shot prompts are used to reinstantiate task contexts in zero-shot settings. The dotted line indicates the ratio between recontextualized zero-shot accuracy and 8-shot accuracy.
Figure S2: Developmental trajectory of task representations over shots. Task vectors are the token activations of the colon token prior to answer generation. We visualize task vectors sourced from the mode best layer across tasks at which task contexts are best restored in a zero-shot setting. Representations for each task are first averaged across samples with the same number of examples in the prompt.
Figure S3: Recontextualized zero-shot accuracy from different format tokens in the prompt in different tasks. The colors indicate different model sizes: blue=GEMMA_V3_4B_PT, yellow=GEMMA_V3_12B_PT, red=GEMMA_V3_27B_PT.
Figure S4: The development of different task representations over layers and context. A. Recontextualized accuracy when each token representation is used to restore task context in zero-shot settings. B. Task identity decoding accuracy across all 14 simple/short-answer tasks in Table 1. C. Task identity decoding accuracy across 9 list operation tasks with shared vocabulary. This subset of tasks includes: choose_first_of_5, choose_middle_of_5, choose_last_of_5, choose_first_last_of_3, choose_first_middle_last_of_5, reverse_all_of_3, reverse_all_of_5, shift_all_of_3, shift_all_of_5.
Figure S5: T-SNE projection of all task vectors across models and tasks.
Appendix C Results on Qwen3 models
Figure S6: Recontextualization accuracy from Qwen3 models for all tasks. Visualization as in Figure S1.
Figure S7: In Qwen3 models, transferrable task representations also condense evidence and activate sporadically at specific tokens. Visualization as in Figures 3A, B and Figure 1C. The hard-to-transfer tasks in Qwen3 models include english-french, choose_first_of_5, count_color_v_animal_3, and count_fruit_v_animal_3. These are the tasks where task vectors extracted from 32-shot prompts failed to recover more than 50% of the 32-shot accuracy on 0-shot prompts, across all model sizes.
Figure S8: Variance and magnitude changes across task vectors extracted following different number of examples. Visualization as in Figure 4A.
Figure S9: T-SNE task representations across all models and tasks.
Figure S10: Developmental trajectory of task representations over shots. Visualization as in Figure S2.
Figure S11: Recontextualized zero-shot accuracy from different format tokens in the prompt. The colors indicate different model sizes: blue=QWEN3_4B, yellow=QWEN3_8B, red=QWEN3_14B.
Figure S12: Task identity decoding accroacy across all 14 simple tasks in Table 1 (A) and 9 list operation tasks with shared vocabulary (B). See Figure S4 caption for a complete list of tasks.
Figure S13: In Qwen3, restored task contexts in longer- and mixed-generation tasks also decay over more output units during generation. Visualization as in Figure 5.